OpenStudy (anonymous):

The length of a rectangle is 5yd longer than its width. If the perimeter of the rectangle is 38yd, find its area.

OpenStudy (bgrg007):

A=12

OpenStudy (anonymous):

@BGrg007 your answer was wrong..

OpenStudy (bgrg007):

ahh sorry about that. that's the length. Area is 35 lol

OpenStudy (anonymous):

wrong again @BGrg007

OpenStudy (anonymous):

A = 84

OpenStudy (nurali):

Let x = width of the rectangle x+5 = length of the rectangle Perimeter = 2(W) + 2(L) = 38 2(x) + 2(x+5) = 38 2x + 2x + 10 = 38 4x + 10 = 38 4x = 28 x= 7 yd Width x+5 = 12 yd Length Area = L*W = 12yd * 7 yd = ----------sq cm.

OpenStudy (anonymous):

84 was right!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

2w+2(w+5)=38 2w+5=19 w=7 yd L=12 area=12 x 7=84 yd^2

OpenStudy (bgrg007):

i was doing it in my head. lol will use paper next time. sorry about the wrong info..twice *_*